The vapour pressure of ethanol (C2H3OH) and 1-propanol (C3H7OH) at 35°C are



100 mmHg and 37.6 mmhg, respectively. Assuming ideal behaviour, calculate the



partial vapour pressures of ethanol and 1-propanol over a solution, in which the



mole fraction of ethanol is 0.3.

1
Expert's answer
2021-12-06T21:43:01-0500

For ethanol partial pressure = mole fraction * Vapour pressure = 0.3 * 100 = 30 mmHg now for propanol partial pressure = 0.7*37.6 = 26.32 mmHg